OpenStudy (anonymous):

What percent of 20 is 18?

OpenStudy (anonymous):

do 20/18 and mulitply by 100

OpenStudy (anonymous):

18/20 = 9/10. 9/10 = .9 .9(100) = 90% 90%

OpenStudy (anonymous):

***18/20